payton:

What is the value of n in the equation –negative StartFraction one-half EndFraction left-parenthesis 2 n plus 4 right-parenthesis plus 6 equals negative 9 plus 4 left parenthesis 2 n plus 1 right-parenthesis.(2n + 4) + 6 = –9 + 4(2n + 1)? n =

darkknight:

Use the order of operations, you need to combine like terms and distribute. Then get x to one side and solve for it. Let me know if and when you are stuck

payton:

im confused

darkknight:

ok, sorry I was afk. So let me walk you through the little steps. (2n + 4) + 6 = –9 + 4(2n + 1) solving for n distribute first: (2n+4)+6 = -9 +(8n+4) combine like terms: 2n+10 = 8n-5 subtract 2n and add 5 on both sides.: 2n - 2n + 10 -5 = 8n - 2n -5 +5 (just subtract 2n from both sides then add 5 on both sides) you get 5 = 6n. divide by 6 to find n

payton:

its okay i was just going to wait anyways

darkknight:

okay, do you understand everything I said above?

darkknight:

You still there lol

payton:

yeah im back

darkknight:

Do you understand everything stated above?

payton:

not really if imma be honest with you

darkknight:

let me rewrite it (2n + 4) + 6 = –9 + 4(2n + 1) solving for n distribute first: (2n+4)+6 = -9 +(8n+4) combine like terms: 2n+10 = 8n-5 subtract 2n : 2n - 2n + 10 = 8n - 2n -5 add 5 on both sides now: 10 + 5 = 6n -5 + 5 15 = 6n you get 15 = 6n. divide by 6 to find n

payton:

im back i went to get a snack im sorry

darkknight:

Its good. Your problem has been solved.

darkknight:

I am the darkknight

payton:

lol do u always say that?

payton:

2.5

darkknight:

ye

payton:

it says thats not it